A recent study examined hearing loss data for 1822 U.S. teenagers. In this sample, 343 were found to have some level of hearing loss. News of this study spread quickly, with many news articles blaming the prevalence of hearing loss on the higher use of ear buds by teens. At MSNBC.com (8/17/2010), Carla Johnson summarized the study with the headline: 1 in 5 U.S. teens has hearing loss, study says. To investigate whether this is an appropriate or a misleading headline, you will conduct a test of significance with the following hypotheses: Null: п-0.20 Alternative : п # 0.20 Use the Theory-Based Inference applet to determine a p-value. Round your answer to 4 decimal places, e.g. 0.7534. Answer *1: the absolute tolerance is/-0.0002 Based on your p-value, there is strong evidence that the proportion of all U.S. teens with some hearing loss is different than 1 in 5 (or 20%) True С False

Using the applet, find a 95% confidence interval for the proportion of US, teens that have some hearing loss. Round your answer to 4 decimal places, e.g. 0.7534 The 95% confidence interval is L Answer *1: the absolute tolerance is +/-0.0005 Answer *2: the absolute tolerance is +/-0.0005 1-2, What is the margin of error for your confidence interval from part (d) of this question? Round your answer to 4 decimal places, eg. 0.7534. Margin of error Answer *1: the absolute tolerance is +/-0.0005 Based on your confidence interval, 0.20 is a plausible value for the proportion of the population that has some hearing loss. True С False Based on your p-value, 0.20 is a plausible value for the proportion of the population that has some hearing loss. True С False Softwar
Answer #1

Ans:

1)Test statistic:

z=(0.1883-0.2)/SQRT(0.2*(1-0.2)/1822)

z=-1.253

p-value=2*P(z<-1.253)=0.2102

2)As,p-value>0.05,we fail to reject H0.

There is not sufficient evidence that proportion is different from 0.2.

False

3)

95% confidence interval for p

=0.1883+/-1.96*sqrt(0.1883*(1-0.1883)/1822)

=0.1883+/-0.0180

=(0.1703, 0.2062)

4)

Margin of error=0.0180

5)

True,0.20 is the plausible value,as 0.20 is included within the interval.

6)

True.